Compute the value of z when ƥ = 0.46, q = 0.54, n = 50 and p0 = 0.50.



1
Expert's answer
2022-05-16T17:42:55-0400
"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.46-0.5}{\sqrt{\dfrac{0.5(1-0.5)}{50}}}"

"=-0.5657"